OpenStudy (anonymous):

What is the equation of the circle with the center (-6, 7) that passes through the point (4, -2)? @flight98

Nnesha (nnesha):

\[\huge\rm (x-h)^2 + (y-k)^2 = r^2\] where (h,k) is the center and passing trough the point (x,y) so replace variables by their values and then solve for r(radius)

Nnesha (nnesha):

(h,k) ( -6,7) center (x,y)(4,-2)

OpenStudy (anonymous):

I'm not sure how to do this....

Nnesha (nnesha):

use that equation i gave u first of all replace h and k by (-6,7) and x ,y by (4,-2)

Nnesha (nnesha):

\[\huge\rm (\color{blue}{x}-\color{red}{h})^2 + (\color{blue}{y}-\color{red}{k})^2 = r^2\] \[\huge\rm (\color{red}{-6,7})\] \[\huge\rm (\color{blue}{4,-2})\]

OpenStudy (anonymous):

These are my options...

Nnesha (nnesha):

okay so plug in (h,k) and (x,y) values into the equation

Nnesha (nnesha):

\[\huge\rm (\color{blue}{4}-\color{red}{h})^2 + (\color{blue}{-2}-\color{red}{k})^2 = r^2\] \[\huge\rm (\color{red}{-6,7})\] \[\huge\rm (\color{blue}{4,-2})\] that's how you should replace x and y by their value now plug in (h,k) values into the equation

OpenStudy (anonymous):

I don't know

Nnesha (nnesha):

i'm not saying tell me the answer just replace h by -6 and k by 7 and then show ur work that's it then we will solve it together